OpenStudy (sette):

Algebra 1 Honors: Dividing Rational Expressions What is the quotient of (a+b/2b^2)/(a^2+3a-18/b^3) ?

OpenStudy (anonymous):

\[\huge\frac{\frac{a+b}{2b^2}}{\frac{a^2+3a-18}{b^3}}\] Is this how your problem appears?

OpenStudy (anonymous):

First, let's multiply by the reciprocal

OpenStudy (anonymous):

\[\frac{a+b}{2b^2}*\frac{b^3}{a^2+3a-18}\]

OpenStudy (anonymous):

\[\huge \frac{a+b}{2}*\frac{b}{a^2+3a-18}\] cancelled the b's

OpenStudy (anonymous):

wow you are so large...

OpenStudy (sette):

My Bad its a+6 /2

OpenStudy (anonymous):

\[\frac{b^2+ab}{2a^2+6a-36}\]

OpenStudy (sette):

Im Sorry.

OpenStudy (anonymous):

dammit

OpenStudy (anonymous):

\[\huge \frac{a+6}{2}*\frac{b}{(a+6)(a-3)}\]

OpenStudy (anonymous):

\[\huge \frac{b}{2(a-3)}\]

OpenStudy (sette):

Thank You

OpenStudy (anonymous):

yw